OpenStudy (tia_sweetheart17):

solve for y. 4x-9y=16

OpenStudy (usukidoll):

-4x on both sides and divide by -9

OpenStudy (anonymous):

Simplifying 4x + -9y = 16 Solving 4x + -9y = 16 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '9y' to each side of the equation. 4x + -9y + 9y = 16 + 9y Combine like terms: -9y + 9y = 0 4x + 0 = 16 + 9y 4x = 16 + 9y Divide each side by '4'. x = 4 + 2.25y Simplifying x = 4 + 2.25y